### v1

@@ 3.2 Prior Specification
To address the structured spatial random effect within a Bayesian hierarchical framework, we assume an ICAR prior for the random vector $\bm{\psi}$. Specifically, the spatially structured random effect is modelled as follows:
@@
where $\bm{W}$ is a binary symmetric $20\times 20$ matrix known as the “weights matrix” or “adjacency matrix” (see, e.g., Banerjee et al.,, 2014; Haining and Li,, 2020). The entries of $\bm{W}$ are defined as follows: $w_{ij}=1$ if regions $i$ and $j$ are neighbours, and $w_{ij}=0$ otherwise. Diagonal elements are set to zero since a region is not considered a neighbour of itself. The diagonal matrix $\bm{D}$ is defined as $\bm{D}=\textbf{Diag}(\bm{W}\cdot\bm{1})$, where Diag represents the diagonal matrix operator, $\bm{1}$ denotes a column vector of ones and $\cdot$ denotes the dot product. The diagonal entries $D_{ii}$ represent the number of neighbours for each region $i$. The scalar $\tau_{\psi}$ is a precision parameter that scales the degree of spatial smoothing in the model: higher values of $\tau_{\psi}$ lead to stronger spatial smoothing, while lower values allow for greater heterogeneity across regions. The full-conditional distribution of the $i$-th spatial random effects $\psi_{i}$ is given by:
@@
where $j\sim i$ indicates that regions $i$ and $j$ are neighbours. As shown in Equation (1), the conditional mean of $\psi_{i}$ is the average of the values of its neighbours, while its variance decreases as the number of neighbours increases. This formulation implies complete spatial dependence, which means that the value of each region is entirely determined by the values of its neighbours (Rue and Held,, 2005).
For what concerns the structured time random effects, following Knorr-Held, (2000) and Berloco et al., (2023), we model the temporal random vector $\bm{\alpha}$ using a first-order auto-regressive (AR(1)) process with the following structure:
@@
where $\tau_{\alpha}$ is the precision parameter and $\rho$ is the auto-regressive coefficient. We assign a uniform prior to $\rho$:
$\rho\sim Beta(1,1),$
such that $\rho$ is restricted to the interval $[0,1]$, reflecting a prior belief in positive autocorrelation among the temporal random effects while preventing explosive dynamics in the AR(1) process.
The gender-specific intercepts are modelled in an additive way. First, we recall that the individual $\psi_{i}$’s cannot be uniquely determined. This issue is resolved by imposing a sum to zero constraint $\sum_{i=1}^{n}\psi_{i}=0$ (Thomas et al.,, 2014) and by modelling the global intercept term $\beta_{0}$ as suggested by Besag and Kooperberg, (1995):
@@
indicating a flat, improper prior for $\beta_{0}$ over the entire real line. A gender-specific shift is introduced through the binary indicator variable $d_{s}$, where $d_{s}=0$ for females and $d_{s}=1$ for males. The gender effect coefficient $\gamma$ allows the definition of gender-specific intercepts as
@@
For the structured spatial and temporal random effects, we assign a Gamma prior to the respective precision parameters. Following the recommendation of Kelsall and Wakefield, (1999), we use a Gamma with shape and rate parameters $\alpha_{\tau}=0.5$ and $\beta_{\tau}=0.0005$ for the spatial precision $\tau_{\psi}$, as it reflects a conservative and weakly informative belief about spatial dependence (Thomas et al.,, 2014). Although originally proposed for spatial effects, we extended this prior to the temporal precision $\tau_{\alpha}$, acknowledging the presence of temporal correlation in our setting, favoring shrinkage while preserving flexibility to capture stronger spatial or temporal structure when supported by the data:
$\tau_{\psi},\tau_{\alpha}$ $\sim\text{Gamma}(\alpha_{\tau},\beta_{\tau}).$
Finally, we assign a prior to the Beta precision parameter $\phi$ following the approach proposed by Figueroa-Zúñiga et al., (2013), which extends the prior structure suggested by Gelman, (2006) specifically for the Beta distribution, by assuming:
@@
To enable variable selection, a prior on the covariate coefficients $\bm{\beta}$ is specified following the SSVS framework, as detailed in Section 3.3.
@@
$\beta_{k}$ Coefficient for the exogenous variables ($k=1,\ldots,p$)
@@ 3.3 Stochastic Search Variable Selection Prior
To model the fixed effects coefficients $\beta_{k},k=1,\dots,p$, we adopt a SSVS prior specification (George and McCulloch,, 1993). This approach combines a spike-and-slab mixture prior to probabilistically determine covariate inclusion. In a nutshell, SSVS enables each regressor to be significantly associated with the response (i.e., different from zero) or shrunk toward zero. This is achieved by assigning a prior that assumes independence between regressors, with marginals defined as a mixture of two Normal distributions, both centered at zero. The spike component corresponds to negligible or irrelevant effects, while the slab component allows for greater variance and captures substantial or practically meaningful associations. Specifically, we assign the following hierarchical prior structure:
$\beta_{k}\mid\sigma^{2}_{k}$ $\sim\mathcal{N}(0,\sigma_{k}^{2}),$
@@
In this formulation, $\omega_{k}$ is a binary latent variable indicating whether the $k$-th covariate is “practically significant” ($\omega_{k}=1$) or not. The prior probability of inclusion of the $k$-th covariate is denoted by $\theta_{k}$, to which we assign a Uniform hyper-prior, avoiding imposing strong prior assumptions about the relevance of individual predictors. The parameters $\tau^{2}$ and $c^{2}\tau^{2}$ represent the variances of the spike and slab components, respectively. The intersection points of the spike and slab densities define a threshold $\zeta$ for practical significance: $\zeta=\pm\tau\sqrt{2c^{2}\log(c)/(c^{2}-1)},$ such that coefficients falling within the interval $\zeta,+\zeta]$ are interpreted as practically negligible. Therefore, in accordance with Rockova et al., (2012), we fix to constant values $\tau$ and $c$ to control the separation between negligible and relevant effects (see Section 3.4). Variable selection is then achieved in practice by either identifying the model with the highest posterior probability, including or excluding covariates based on their marginal posterior inclusion probabilities, or evaluating the extent to which the marginal posterior distributions of the regression coefficients are shrunk toward zero (for further details see Rockova et al.,, 2012).
@@ 3.4 Computational Details
A summary of all the model parameters is reported in Table 2. To conduct Bayesian inference, we implemented the proposed spatio-temporal hierarchical model using the Nimble package (de Valpine et al.,, 2017, 2024), which enables efficient specification and sampling of complex models. The final model configuration was determined after extensive empirical testing to ensure posterior convergence and computational efficiency. To improve convergence of the MCMC algorithm, we normalized the variables using min-max scaling and standardization, as appropriate. For what concerns the SSVS specification, we set the slab scaling factor to $c=4000$, and the intersection term to $\zeta=0.001$, reflecting the threshold of practical significance we determined through iterative exploration and consideration of the magnitude of the coefficients. We derived the standard deviation of the spike using the relationship: $\tau={\zeta}/{\sqrt{{2\log(c)c^{2}}/(c^{2}-1)}},$ which yielded $\tau=0.00025$, a value corresponding to a very small standard deviation for the spike around 0. The slab variance was then calculated as $\tau^{2}c^{2}$, resulting in $\tau^{2}c^{2}\approx 1$. This configuration aligns the slab variance with the suggested scale of variation for coefficients under the logit link function, as suggested by McElreath, (2016).
We performed parameter estimation via the MCMC algorithm in Nimble, with 150,000 iterations across two chains, a burn-in period of 50,000 iterations and a thinning interval of 10, yielding a 10,000 posterior sample size.
@@ 4 Results
Parameter (Gender) Mean SD CV Skewness Kurtosis
@@
Figure 1: Posterior distributions of the gender-specific intercepts: $\xi_{1}$ for females (in pink) and $\xi_{2}$ for males (in light blue).
@@
Figure 1 depicts the posterior density estimates of the gender-specific intercepts. The clear separation between the two distributions emphasizes the significant role gender plays in modeling obesity. This result highlights the substantial contribution of gender-specific effects in explaining differences in obesity prevalence.
@@
Figure 2: Posterior density estimation of the space random effects $\psi_{i}$, coloured by geographical zone (NO=North West, NE=North East, C=Centre, S=South, SI=Islands, see Table 1).
@@
Figure 3: Posterior mean values of the spatial random effects $\psi_{i}$ for each Italian region.
@@
The structured spatial random effects can be visualized through the posterior densities plotted in Figure 2. The plot shows the posterior densities for all regions, highlighting the contribution of regional random effects to the overall obesity estimates according to the model. Overall, the ICAR prior effectively induces a spatial structure that captures regional dependencies while allowing for evident heterogeneity, as reflected in the varying magnitudes of $\psi_{i}$ across macro areas and individual regions. To complement the density plots, Figure 3 provides a map representation of the posterior mean values of the spatial random effects, facilitating the interpretation of their spatial distribution. While spatial proximity generally corresponds to similar values of the spatial random effect, notable exceptions suggest that additional determinants beyond geography influence regional obesity rates. As shown by the posterior summary statistics in Table 8 of Appendix 11, regions with negative posterior mean values, such as Piemonte (-0.61), Liguria (-0.62), Trentino-Alto Adige (-0.49), and Sardegna (-0.71, the lowest), are associated with lower obesity rates. In contrast, with the exception of Calabria, all southern regions, particularly Molise (0.89, the highest), as well as Umbria (0.29), exhibit strong positive spatial effects. The uncertainty in spatial effects varies across regions, as reflected in the standard deviations of the posterior distributions. Trentino-Alto Adige and Calabria (SD = 0.12, the highest), Campania, and Sicilia exhibit greater variability, whereas Veneto, Friuli-Venezia Giulia, Umbria, Marche, Lazio and Abruzzo show more spiked posterior estimates, with Abruzzo having the lowest SD (0.06). Additionally, the distributions reveal a skewness pattern: northern regions such as Piemonte, Lombardia, Liguria, and Lazio exhibit right-skewness, while southern regions generally show left-skewness. Notably, all the regions have kurtosis values exceeding 3, suggesting pronounced posterior tails, with Puglia being the region with the heaviest tails. To summarize, while spatial correlation in obesity rates is evident, the variation across regions is not as smooth as one might expect from a prior ICAR specification. Some regional posterior densities deviate strongly from those of neighbouring regions, suggesting localized spatial heterogeneity that is not entirely captured by the prior smoothing assumptions. As regards the posterior summary statistics for the spatial precision parameter $\tau_{\psi}$, we refer the reader to Appendix 11.
@@
Figure 4: Posterior density estimation of the time random effects $\alpha_{t}$, coloured by year.
@@
Table 4: Posterior probability that the temporal effect $\alpha_{t}$ exceeds zero (PP0).
@@
[…]
As confirmed by the posterior summary statistics shown in Table 9 of Appendix 11, the mean values of the $\alpha_{t}$ parameters range from negative values, such as $\alpha_{2}$ and $\alpha_{3}$, which have the lower mean (-0.08), to positive values, with $\alpha_{12}$ exhibiting the highest one (0.41). These differences highlight the varying influence of the time random effects across different time points. While some periods exhibit negative random effects, indicating a lower-than-expected influence on the obesity rates, others demonstrate positive effects, which may indicate periods of higher-than-expected influence. The standard deviations of the $\alpha_{t}$ parameters are relatively stable, ranging from 0.12 to 0.15. This consistency suggests that while the magnitude of time random effects may differ across time points, the variability within each time point is relatively homogeneous. With few exceptions, the $\alpha_{t}$ coefficients show a positive skewness, consistent in the second half of the period, indicating again the increase in the temporal effect. All of the $\alpha_{t}$ parameters exhibit kurtosis values greater than 3, interestingly showing the heavy tails of all the posterior distributions of the time random effect.
@@
[…]
We briefly mention that the auto-regressive parameter $\rho$ show a negatively skewed posterior distribution, with a posterior mean of 0.85, suggesting quite strong temporal dependence from year to year (see Table 5). As regards the posterior summary statistics for the temporal precision parameter $\tau_{\alpha}$, we refer the reader to Table 10 in Appendix 11.
@@
The posterior summary statistics of the Beta precision parameter $\phi$ reported in Table 6 are not straightforward to interpret, primarily because they depend on the scale of the data. The parameter becomes more interesting when compared between models: a model with a higher posterior precision parameter suggests a better ability to explain the variability of the Beta-distributed response variable, due to the relation that links the variance to the precision $\phi$ (see Section 3). We just highlight that the Bayesian updating process was very pronounced, as the prior distribution, which spans a wide range (between 0 and $2500$), is effectively updated to a more concentrated range around the mean of 37.67. The reduction in uncertainty is very high, as reflected by the low coefficient of variation equal to 0.06. Finally, we observe only a mild positive skewness and a kurtosis value nearly identical to that of a Normal distribution.
@@
Figure 5: Posterior probability evaluation of the inclusion parameters $\omega_{k}$ for all the covariates (for the description of the covariates see Table 7 in Appendix 10).
@@
As discussed in Section 3.3, we adopt the SSVS approach to perform inference on the fixed-effect coefficients $\beta_{k}$. Consequently, the evaluation of their posterior distribution is intrinsically linked to the variable selection induced by the spike and slab prior. The SSVS hierarchical prior specification can make inference on the $\beta_{k}$ coefficients challenging, particularly when their posterior distributions are bi-modal, with samples alternating between values near zero and substantially different from zero. This complicates the inferential process. However, we point out that the primary objective of this work is to investigate the spatio-temporal dynamics of obesity rates across Italian regions, rather than exploring or assessing causal relationships between the covariates and the response variable. The inclusion of these covariates is primarily to control for any additional effect that may arise along with the random effects, which are of greater interest in this study. Thus we do not conduct detailed posterior evaluation of the $\beta_{k}$ coefficients. Nevertheless, given the valuable insights offered by the SSVS approach, we perform a brief analysis of the inclusion probabilities to assess which covariates are associated with obesity, after accounting for spatial and temporal random effects, as well as gender. Specifically, the $\omega_{k}$ parameter plays a central role in determining whether the $k$-th covariate should be included in the model, based on the probability parameter $\theta_{k}$ (see Section 3.3). By counting the proportion of MCMC iterations in which $\omega_{k}=1$, we obtain an empirical inclusion probability for each covariate. Figure 5 presents these posterior inclusion probabilities. The only covariates that are consistently included in the model are life expectancy, life expectancy in good health and overweight. All the remaining covariates exhibit very low posterior inclusion probabilities. This suggests that, once the random effects and gender-specific intercepts are incorporated, only three covariates need to be included according to our model specification. The other covariates do not prove to be related with obesity rates and only add unnecessary complexity to the model.

### v2

@@ 3.2 Prior specification
For the spatially structured random effect $\bm{\psi}$ we assume an ICAR prior, a standard and computationally convenient specification for structured areal effects, as it encourages local smoothing by borrowing strength across neighboring regions [33, 2]. Specifically, the spatially structured random effect is modelled as follows:
@@
where $\bm{W}$ is a binary symmetric $20\times 20$ matrix known as the “weights matrix” or “adjacency matrix” [see, e.g., 2, 20]. The entries of $\bm{W}$ are defined as follows: $w_{ij}=1$ if regions $i$ and $j$ share a common boundary (i.e., are first–order contiguous neighbours) and $w_{ij}=0$ otherwise. Diagonal elements are set to zero since a region is not considered a neighbour of itself. The diagonal matrix $\bm{D}$ is defined as $\bm{D}=\textbf{Diag}(\bm{W}\cdot\bm{1})$, where Diag represents the diagonal matrix operator, $\bm{1}$ denotes a column vector of ones and $\cdot$ denotes the dot product. The diagonal entries $D_{ii}$ represent the number of neighbours for each region $i$. The scalar $\tau_{\psi}$ is a precision parameter that scales the degree of spatial smoothing in the model: higher values of $\tau_{\psi}$ lead to stronger spatial smoothing, while lower values allow for greater heterogeneity across regions. The full-conditional distribution of the $i$-th spatial random effects $\psi_{i}$ is given by:
@@
where $j\sim i$ indicates that regions $i$ and $j$ are neighbours. As shown in Equation (1), the conditional mean of $\psi_{i}$ is the average of the values of its neighbours, while its variance decreases as the number of neighbours increases.
For what concerns the structured time random effects, following Knorr-Held [25] and Berloco et al. [4], we model the temporal random vector $\bm{\alpha}$ using a first-order auto-regressive (AR(1)) process with the following structure:
@@
where $\tau_{\alpha}$ is the precision parameter and $\rho$ is the auto-regressive coefficient. We assign a uniform prior to $\rho$ over the interval $(0,1)$:
$\rho\sim\text{Beta}(1,1),$
reflecting a prior belief in positive autocorrelation among the temporal random effects while preventing explosive dynamics in the AR(1) process.
The gender-specific intercepts are modelled in an additive way. First, we recall that the individual $\psi_{i}$’s cannot be uniquely determined. This issue is resolved by imposing a sum to zero constraint $\sum_{i=1}^{n}\psi_{i}=0$ [35] and by modelling the global intercept term $\beta_{0}$ as suggested by [5]:
@@
indicating a flat, improper prior for $\beta_{0}$ over the entire real line. A gender-specific shift is introduced through the binary indicator variable $d_{s}$. The gender effect coefficient $\gamma$ allows the definition of gender-specific intercepts as
@@
For the structured spatial and temporal random effects, we assign a Gamma prior to the respective precision parameters $\tau_{\psi}$ and $\tau_{\alpha}$. In particular, following the recommendation of Kelsall and Wakefield [24], for the spatial precision $\tau_{\psi}$ we use a $\text{Gamma}(\alpha_{\tau},\beta_{\tau})$ prior with $\alpha_{\tau}=0.5$ and $\beta_{\tau}=0.0005$; this reflects a conservative and weakly informative belief about spatial dependence [35]. Although originally proposed for spatial effects, we extended this prior to the temporal precision $\tau_{\alpha}$, acknowledging the presence of temporal correlation in our setting, favoring shrinkage while preserving flexibility to capture stronger temporal structure when supported by the data.
Finally, we assign a prior to the Beta precision parameter $\phi$ following the approach proposed by Figueroa-Zúñiga et al. [13], which extends the prior structure suggested by Gelman [15] specifically for the Beta distribution, by assuming:
@@
To complete the Bayesian hierarchical formulation, we define the prior for the covariate coefficients $\bm{\beta}$ using the SSVS approach, as discussed in Section 3.3.
@@
$\beta_{k}$ Regressor coefficient ($k=1,\ldots,p$)
@@ 3.3 Stochastic Search Variable Selection Prior
To model the fixed effects coefficients $\beta_{k},k=1,\dots,p$, we adopt a SSVS prior specification [17]. This approach combines a spike-and-slab mixture prior to probabilistically determine covariate inclusion. In a nutshell, SSVS enables each regressor to be significantly associated with the response (i.e., coefficient different from zero) or shrunk toward zero. This is achieved by assigning a prior that assumes independence between regressors, with marginals defined as a mixture of two Normal distributions, both centered at zero. The spike component corresponds to negligible or irrelevant effects, while the slab component allows for greater variance and captures substantial or practically meaningful associations. Specifically, we assign the following hierarchical prior structure:
$\beta_{k}\mid\sigma^{2}_{k}$ $\sim\mathcal{N}(0,\sigma_{k}^{2}),\qquad k=1,\ldots,p$
@@
In this formulation, $\omega_{k}$ is a binary latent variable indicating whether the $k$-th covariate is “practically significant” ($\omega_{k}=1$) or not. The prior probability of inclusion of the $k$-th covariate is denoted by $\theta_{k}$, to which we assign a Uniform hyper-prior, avoiding imposing strong prior assumptions about the relevance of individual predictors. The parameters $\tau^{2}$ and $c^{2}\tau^{2}$ represent the variances of the spike and slab components, respectively. The intersection points of the spike and slab densities define a threshold $\zeta$ for practical significance: $\zeta=\pm\tau\sqrt{2c^{2}\log(c)/(c^{2}-1)},$ such that coefficients falling within the interval $\zeta,+\zeta]$ are interpreted as practically negligible. Therefore, in accordance with Rockova et al. [32], we fix $\tau$ and $c$ to constant values in order to control the separation between negligible and relevant effects (see Section 3.4). Variable selection is then achieved in practice by either identifying the model with the highest posterior probability, including or excluding covariates based on their marginal posterior inclusion probabilities, or evaluating the extent to which the marginal posterior distributions of the regression coefficients are shrunk toward zero [for further details see 32].
@@ 3.4 Computational details
A summary of all the model parameters is reported in Table 1. To conduct Bayesian inference, we implemented the proposed spatio-temporal hierarchical model using the Nimble package [36, 37], which enables efficient specification and sampling of complex models. To improve convergence of the MCMC algorithm, we normalized the variables using min-max scaling and standardization, as appropriate.
With regard to the SSVS specification, we set the slab scaling factor to $c=4000$, and the intersection term to $\zeta=0.001$, reflecting the threshold of practical significance identified through iterative exploration and consideration of the magnitude of the coefficients (we refer to Appendix E for further details about the choice and sensitivity analysis of SSVS parameters). This resulted in $\tau=0.00025$ (see Section 3.3), a value corresponding to a very small standard deviation for the spike around 0. The slab variance, given by $\tau^{2}c^{2}$, is approximately equal to 1. This configuration aligns the slab variance with the suggested scale of variation for coefficients under the logit link function, as suggested by McElreath [27].
We ran two parallel chains with overdispersed initial values for a total of 150,000 iterations each. We discarded an initial burn-in period of 50,000 iterations from each chain to ensure convergence to the stationary distribution. To reduce storage requirements and mitigate the effects of residual autocorrelation in the MCMC output, we thinned posterior samples by retaining every 10th iteration. This resulted in a total of 10,000 posterior draws used for inference. The total computation time required to fit the model was approximately 30 minutes per MCMC chain on a standard laptop.
We assessed convergence and sampling efficiency using multiple diagnostics, including rank-normalized $\widehat{R}$ statistics and effective sample size (ESS) measures as proposed by Vehtari et al. [39], together with the more classic ESS and $\widehat{R}$ [31, 16, 6]. Detailed convergence diagnostics are reported in Appendix F.
@@ 4 Results
Parameter Mean SD CV Skewness Kurtosis
@@
Figure 2: Posterior distributions of the gender-specific intercepts: $\xi_{1}$ for females (in pink) and $\xi_{2}$ for males (in light blue)
@@
Figure 2 depicts the posterior density estimates of the gender-specific intercepts. The clear separation between the two distributions highlights the substantial contribution of gender-specific effects in explaining differences in obesity prevalence.
@@
Figure 3: Posterior density of the space random effects $\psi_{i}$, coloured by geographical zone (NO=North West, NE=North East, C=Centre, S=South, SI=Islands, see Table 6)
@@
Figure 4: Posterior mean values of the spatial random effects $\psi_{i}$ for each Italian region
@@
The structured spatial random effects can be visualized through the posterior densities plotted in Figure 3. The plot shows the posterior densities for all regions, highlighting the contribution of regional random effects to the overall obesity estimates according to the model. Overall, the ICAR prior effectively induces a spatial structure that captures regional dependencies while allowing for evident heterogeneity, as reflected in the varying magnitudes of $\psi_{i}$ across macro areas and individual regions. To complement the density plots, Figure 4 provides a map representation of the posterior mean values of the spatial random effects, facilitating the interpretation of their spatial distribution. While spatial proximity generally corresponds to similar values of the spatial random effect, notable exceptions suggest that additional determinants beyond geography influence regional obesity rates. As shown by the posterior summary statistics in Table 8 of Appendix C, regions with negative posterior mean values, such as Piemonte (-0.61), Liguria (-0.62), Trentino-Alto Adige (-0.49), and Sardegna (-0.71, the lowest), are associated with lower obesity rates. In contrast, with the exception of Calabria, all southern regions, particularly Molise (0.89, the highest), as well as Umbria (0.29), exhibit strongly positive spatial effects. The uncertainty in spatial effects varies across regions, as reflected in the standard deviations (SD) of the posterior distributions. Trentino-Alto Adige and Calabria (SD = 0.12, the highest), Campania, and Sicilia exhibit greater variability, whereas Veneto, Friuli-Venezia Giulia, Umbria, Marche, Lazio and Abruzzo show more spiked posterior estimates, with Abruzzo having the lowest SD (0.06). Additionally, the distributions reveal a skewness pattern: northern regions such as Piemonte, Lombardia, Liguria, and Lazio exhibit right-skewness, while southern regions generally show left-skewness. Notably, all the regions have kurtosis values exceeding 3, suggesting pronounced posterior tails, with Puglia being the region with the heaviest tails. To summarize, while spatial correlation in obesity rates is evident, the variation across regions is not as smooth as one might expect from a prior ICAR specification. Some regional posterior densities deviate strongly from those of neighbouring regions, suggesting localized spatial heterogeneity that is not entirely captured by the prior smoothing assumptions. As regards the posterior summary statistics for the spatial precision parameter $\tau_{\psi}$, we refer the reader to Appendix C.
@@
Figure 5: Posterior density of the time random effects $\alpha_{t}$, coloured by year
@@
Table 3: Posterior probability that the temporal effect $\alpha_{t}$ exceed zero (PP0).
@@
As confirmed by the posterior summary statistics shown in Table 9 of Appendix C, the mean values of the $\alpha_{t}$ parameters range from negative values, such as $\alpha_{2}$ and $\alpha_{3}$, which have the lowest mean (-0.08), to positive values, with $\alpha_{12}$ exhibiting the highest one (0.41). These differences highlight the varying influence of the time random effects across different time points. While some periods exhibit negative random effects, indicating a lower-than-expected influence on the obesity rates, others demonstrate positive effects, which may indicate periods of higher-than-expected influence. The standard deviations of the $\alpha_{t}$ parameters are relatively stable, ranging from 0.12 to 0.15. This consistency suggests that while the magnitude of time random effects may differ across time points, the variability within each time point is relatively homogeneous.
@@
The autoregressive parameter $\rho$ exhibits a negatively skewed posterior distribution, with a posterior mean of 0.85 (Table 4), indicating strong temporal persistence in regional obesity rates. This suggests that obesity levels in a given year are highly correlated with those in previous years, resulting in a smooth temporal evolution rather than abrupt year-to-year changes. Such persistence is coherent with the influence of slowly evolving underlying factors, rather than short-term shocks. Although broader socioeconomic, behavioral, or epidemiological processes may have contributed to this upward trend, the aggregated nature of the data and the lack of information on specific policies or interventions prevent a more detailed analysis. We therefore interpret the estimated value of $\rho$ primarily as capturing long-run temporal dependence in obesity dynamics, rather than effects associated with specific policy actions or events. As regards the posterior summary statistics for the temporal precision parameter $\tau_{\alpha}$, we refer the reader to Table 10 in Appendix C.
@@
The posterior summary statistics of the Beta precision parameter $\phi$ reported in Table 5 are not straightforward to interpret, primarily because they depend on the scale of the data. The parameter becomes more interesting when compared between models: a model with a higher posterior precision parameter suggests a better ability to explain the variability of the Beta-distributed response variable, due to the relation that links the variance to the precision $\phi$ (see Section 3). We just highlight that the Bayesian updating process was very pronounced, as the prior distribution, which spans a wide range (between 0 and $2500$), is effectively updated to a more concentrated range around the mean of 37.67. The reduction in uncertainty is very high, as reflected by the low coefficient of variation equal to 0.06.
@@
Figure 6: Posterior probability evaluation of the inclusion parameters $\omega_{k}$ for all the covariates (for the description of the covariates see Table 7 in Appendix B)
@@
As discussed in Section 3.3, we adopt the SSVS approach to perform inference on the fixed-effect coefficients $\beta_{k}$. Consequently, the evaluation of their posterior distribution is intrinsically linked to the variable selection induced by the spike and slab prior. The SSVS hierarchical prior specification can make inference on the $\beta_{k}$ coefficients challenging, particularly when their posterior distributions are bi-modal, with samples alternating between values near zero and substantially different from zero. This complicates the inferential process. However, we point out that the primary objective of this work is to investigate the spatio-temporal dynamics of obesity rates across Italian regions, rather than exploring or assessing causal relationships between the covariates and the response variable. The inclusion of these covariates is primarily to control for any additional effect that may arise along with the random effects, which are of greater interest in this study. Thus we do not conduct detailed posterior evaluation of the $\beta_{k}$ coefficients. Nevertheless, given the valuable insights offered by the SSVS approach, we perform a brief analysis of the inclusion probabilities to assess which covariates are associated with obesity, after accounting for spatial and temporal random effects, as well as gender. Specifically, the $\omega_{k}$ parameter plays a central role in determining whether the $k$-th covariate should be included in the model, based on the probability parameter $\theta_{k}$ (see Section 3.3). By counting the proportion of MCMC iterations in which $\omega_{k}=1$, we obtain an empirical inclusion probability for each covariate. Figure 6 presents these posterior inclusion probabilities. The only covariates that are consistently included in the model are “life expectancy”, “life expectancy in good health” and “overweight” (see Table 7). All the remaining covariates exhibit very low posterior inclusion probabilities. This suggests that once the random effects and gender-specific intercepts are incorporated, only three covariates need to be included according to our model specification. The other covariates do not prove to be related to obesity rates and only add unnecessary complexity to the model. We observe that this is consistent with the structure of the model. In particular, the spatial and temporal random effects act as latent proxies for a wide range of unobserved regional characteristics, such as cultural habits, health system features, environmental conditions, socioeconomic context, or persistent lifestyle patterns. As a consequence, much of the cross-regional variability that might otherwise be captured by observable covariates is absorbed by the spatial and temporal terms, reducing the marginal contribution of many individual predictors. This highlights that, at the regional scale, unobserved or hard-to-measure contextual factors play a dominant role relative to the socioeconomic covariates available in the dataset.
To assess the sensitivity of the SSVS procedure to hyperparameter choices, we conducted a sensitivity analysis to determine whether different SSVS settings yield distinct shrinkage patterns. The full analysis is reported in Appendix E. Briefly, we tested four alternative configurations based on different intersections between the spike and slab densities (see Section 3.3). While posterior inclusion probabilities vary slightly across specifications, the same three covariates consistently emerge as the only ones not shrunk toward zero. All the remaining covariates have posterior inclusion probabilities below the commonly used 0.5 threshold, with a minor exception in one specification (see Appendix E for further details).
4.1 Predictive performance